User name nanda
Log entry time 18:42:32 on November 2,2006
Entry number 188761
Compton photon background rates were well above 1kHz/uA. Too high to see Compton signal. We tweaked the beam orbit and focussing in the chicane but the background reductions was marginal. Subsequent dog leg correctin in the 5th arc provided the magic solution bringing down the rates to around 30Hz/uA with 5uA of beam current.
However, going to higher currents, the background rate shoots upto 20 kHz/uA. The transition happens at about 50 uA. MCC will work further to debug this problem.
For now the Compton tune is good to 50uA.